OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14

OpenStudy (anonymous):

Should we use elimination? Let's multiply the first equation by -2 to get -6x + 4y = -12 6x - 4y = 14 Add the columns and what do you get?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7? -6x + 6x = 0 4y - 4y = 0 -12 + 14 = ?? Seven?

OpenStudy (anonymous):

someone help me

OpenStudy (anonymous):

Did you multiply the first equation by 2? Do you agree that -6x + 6x = 0?

OpenStudy (anonymous):

someone help me Hmmm I helped a few times but no thanks or metals awarded and meverett04 is helping you too